What is the ΔrG for the following reaction (in kJ mol-1) at 298 K?

2 Si (s) + 3 H2 (g) ⇌ Si2H6(g)

The conditions for this reaction are:

PH2 = 1.83 bar

PSi2H6 = 0.96 bar

You will also need to use Appendix II in your textbook (containing standard Gibbs energies of formation).

ΔGf(Si2H6) = 127.3 kJ/mol

ΔGf(H2) = 0 kJ/mol
Answer #1

2 Si (s) + 3 H2 (g) ⇌ Si2H6(g)

PH2 = 1.83 bar

PSi2H6 = 0.96 bar

Qp   = PSi2H6/P^3H2

        = 0.96/(1.83)^3   = 0.1566

2 Si (s) + 3 H2 (g) ⇌ Si2H6(g)

ΔG0rxn = ΔGf products - ΔGf reactants

           = 127.3-(2*0 + 3*0)

            = 127.3KJ/mole

             = 127300J/mole

ΔG = ΔG0rxn + RTlnQp

        = 127300 + 8.314*298ln(0.1566)

        = 127300 + 8.314*298*-1.854

         = 122706J/mole

         = 122.706KJ/mole >>>>answer